# SonicSieve: Bringing Directional Speech Extraction to Smartphones Using Acoustic Microstructures

### v1

@@ 3.1. Speech-aware acoustic microstructure for smartphones
3.1.3. Attaching the microstructure to smartphones.
Finally, we describe how the acoustic microstructure is coupled to the in-line microphone of standard earphones. We note that this design choice allows the microstructure to be physically connected to any device with an audio jack, eliminating the need for custom, device-specific cases or attachments. The final design consists of two main components: the top part houses the optimized microstructure, while the bottom part serves as a holder that attaches to the smart device. A silicone washer is placed at the interface between the two components to minimize sound leakage.
3.2. Directional speech extraction network
As described in Sec. 3.1, our approach encodes direction-specific features in the signal recorded by the microphone inside the microstructure (in-line mic at earphone), while capturing an unmodified reference signal using a co-located external microphone (built-in top mic at smartphone). By calculating the ratio between these two signals (Eq. 3), we can extract direction-specific features when a single sound source is present. Although real-world acoustic environments are considerably more complex with multiple sound sources, this fundamental principle remains effective for extracting spatial features. These features, combined with the raw two-channel recordings, provide our neural network with the necessary information to perform directional speech separation.
To support applications such as audio live streaming via existing communication protocols, real-time processing capability on smartphones is essential. To achieve this, we start our neural network design from a state-of-the-art real-time sound extraction network (Chen et al., 2024) and adapt it to our of angular-based directional speech extraction.
Pre-processing Module. Our system takes a 2-channel audio input represented as $x=[x_{1},x_{2}]^{T}\in R^{2\times T}$, where $x_{1}$ and $x_{2}$ correspond to the reference microphone and in-microstructure microphone signals respectively, with $T$ representing the total audio sample length. To accommodate model lookahead requirements, we reshape the input with zero padding to $x\in R^{2\times(T+\sigma)}$, where $\sigma$ is the lookahead sample count.
We implement a learnable Short-Time Fourier Transform (STFT) encoder that provides a sufficiently large encoder window size, enhancing model stability across different reverberation environments (Cord-Landwehr et al., 2022). After applying the STFT encoder with discrete Fourier transform length $L$ and step size $\delta$, the output for each microphone channel $i$ becomes $X_{i}[f,t]\in R^{F\times\frac{(T+\sigma-L)}{\delta}}$, where $F$ represents the number of frequency bins.
@@ 3.2. Directional speech extraction network
Figure 8. Real-world evaluation setup. (a–e) Five different test environments used for evaluating our system. The room dimensions are as follows (in meters): length $\times$ width $\times$ height. Room A, 6.1 $\times$ 4.6 $\times$ 3.7 (a); Room B, 4.9 $\times$ 3.6 $\times$ 3.5 (b); Room C, 6.1 $\times$ 4.6 $\times$ 3.7 (c); Room D, 7.0 $\times$ 4.9 $\times$ 2.7 (d); Room E, 11.5 $\times$ 9.1 $\times$ 2.9 (e). In each room we play sound samples from an external speaker (model name: JBL Flip 5), and record it with multiple devices on our data capture rig.
@@
Feature Extraction. We developed a specialized feature extraction algorithm to isolate and enhance the direction-specific characteristics encoded by the microstructure. As presented in Eq. 3, the directional signature $M_{\theta}$ can be estimated using the ratio between the in-microstructure and reference microphone signals:
(4) $M_{\theta}\approx\frac{X_{2}[f,t]}{X_{1}[f,t]}$
Specifically, we calculate the phase and magnitude of this ratio, which correspond to the interchannel phase differences (IPD) and interchannel level differences (ILD) — spatial cues extensively studied in binaural hearing research (Blauert, 1997; Gu and Luo, 2024). Mathematically, these are calculated as:
(5) $\text{IPD}[f,t]$ $=\angle X_{2}[f,t]-\angle X_{1}[f,t]$
$\text{ILD}[f,t]$ $=20\log\left\{|X_{2}[f,t]|/|X_{1}[f,t]|\right\}$
where $\angle$ and ||$\cdot$|| ⋅ | denote phase and magnitude calculations of the complex number.
To facilitate neural network learning, we compute the cosine and sine transformations of the IPD, which provide a continuous representation of the circular phase information:
We then normalize each feature across frequency bands to ensure consistent scale and improve training stability. For ILD, we apply per-frequency normalization to zero mean and unit variance, which reduces the impact of frequency-dependent gain variations inherent in the microstructure design. Similarly, while the cosine and sine IPD components are naturally bounded, normalization helps balance their influence across the spectrum where phase relationships have varying reliability at different frequencies.
The complete feature set combines the real and imaginary components of the STFT output from both microphone channels (4 channels) with our three spatial features (cosIPD, sinIPD, and ILD), resulting in 7 total feature channels. We process these through a 2D convolutional block to encode them into a representation $X\in R^{c\times F\times\frac{(T+\sigma-L)}{\delta}}$, where $c$ is the feature channel number expanded by the convolutional layer.
Angle Encoder. We represent the user’s directional selection through a 6-sector encoding system. Selected sectors are marked with 1 in a one-hot vector, while adjacent sectors receive a value of 0.25 to provide smooth spatial context. This vector passes through embedding refinement, fully connected layers, and reshaping operations to create directional embeddings that match our feature dimensions. After LayerNorm, these embeddings ($E(x)$) modulate features in each separation block through Feature-wise Linear Modulation (FiLM).
Real-time Separation Blocks. Our core processing pipeline consists of N sequential LSTM-based blocks optimized for real-time operation. Each block begins with LayerNorm and PReLU activation before using convolutional downsampling to reduce computational demands for the bidirectional LSTM. After frequency-domain processing, a deconvolution layer with skip connection restores the original dimensions. A unidirectional (causal) LSTM with a fully connected layer then processes temporal relationships, maintaining real-time performance. The FiLM layer integrates directional information by modulating features according to E(x): $E_{sep}=X^{\prime}*Conv1D(E(x))+Conv1D(E(x))$, where X’ represents processed features.
The final output undergoes transposed convolution and inverse STFT to generate time-domain audio. For continuous processing, we output T samples for playback while retaining $\sigma$ samples for future chunk processing, enabling smooth real-time operation.
Figure 9. System performance compared to different baselines and by selected sector. (a) Effect of selected sector index on SI-SDR improvement, with and without microstructure. (b) SI-SDR improvement with and without microstructure, as a function of the number of simultaneously selected sectors. (c) Input vs. output SI-SDR of our model across all samples, grouped by 1-sector, 2-sector, and 3-sector selections. Points above the dashed line indicate cases where SI-SDR improved.
[…]
This figure presents three graphs evaluating system performance through SI-SDR (Scale-Invariant Signal-to-Distortion Ratio) measurements across different conditions: (a) A line graph showing the effect of selected sector index (x-axis, numbered 1-6) on SI-SDR improvement in dB (y-axis, ranging 2-7 dB). Two lines compare performance: With microstructure (light green line with circles): Shows significantly higher SI-SDR values (between 3.6-6.8 dB) with a U-shaped pattern where sectors 1 and 6 perform best. The SI-SDR number for each sector is 6.4, 4.7, 4.4, 4.1, 3.6, 6.8. Without microstructure (dark blue line with squares): Shows consistently lower performance (between 1.9-2.8 dB) with a slight downward trend as sector index increases. The SI-SDR number for each sector is 2.7, 2.0, 2.5, 2.4, 2.1, 1.9. (b) A bar graph comparing SI-SDR improvement (y-axis, in dB) across different numbers of simultaneously queried sectors (x-axis, 1-3 sectors). For each number of sectors, two bars compare: With microstructure (green bars): Shows 5.0 dB for 1 sector, 4.2 dB for 2 sectors, and 3.9 dB for 3 sectors. Without microstructure (blue bars): Shows substantially lower values of 2.3 dB, 1.9 dB, and 1.7 dB respectively. (c) A scatter plot showing input vs. output SI-SDR (x-axis: Original SI-SDR ranging from -25 to 15 dB; y-axis: Output SI-SDR ranging from -30 to 20 dB). Points are color-coded by query type: 1 Sector Query (red dots), 2 Sectors Query (blue dots), 3 Sectors Query (orange dots). A diagonal dashed line indicates ”Zero improvement” reference. Most data points cluster above this line (especially in the -5 to 5 dB original SI-SDR range), demonstrating positive enhancement, with the upper-right quadrant labeled ”Enhancement” in green.
@@ 3.3. Real-world dataset
To ensure our system works well in different real-world scenarios, we augmented existing audio datasets by replaying them through a speaker in five different rooms, including meeting rooms, conference rooms and classrooms (Fig. 8).
@@
Data sources for real-world data augmentation. We use the VCTK (Veaux et al., 2017) and LibriTTS (Zen et al., 2019) datasets which contains clips of human speech. All audio clips are resampled to 24 kHz to balance audio quality and real-time requirements.
@@
Dataset collection setup. Our data collection setup includes a JBL speaker [^1] https://www.jbl.com/bluetooth-speakers/JBL+FLIP+5-.html that plays sounds, and a receiver rig with two recording devices: a microphone array from MiniDSP [^2] https://www.minidsp.com/products/usb-audio-interface/uma-8-microphone-array and a Motorola Edge smartphone with a microstructure. For our experiments, we use the smartphone’s in-line mic design with its top microphone as the reference. The receiver rig also features six LED lasers that divide the room into sectors to guide precise placement of the speaker.
To collect data, we placed the receiver in 3 random positions within each room. For each receiver spot, we positioned the transmitter in 5 different locations across all 6 sectors. At each specific transmitter-receiver pair position configuration, 6 random speech clips were played over the air. This process was repeated for 5 different rooms, yielding a total dataset of 2700 clips (6 clips × 6 sectors × 5 transmitter positions × 3 receiver positions × 5 rooms).
@@
In software, we generated sound mixtures based on the number of target sectors. For single-target scenarios, we created 6 possible sector combinations; for two-target scenarios, 15 combinations; and for three-target scenarios, 20 combinations. This resulted in 1200 sound mixtures for each transmitter-receiver position configuration per room in single target sector scenarios, 900 mixtures for two-target sectors, and 900 mixtures for three-target sectors. We controlled the Signal-to-Noise Ratio (SNR) of signals in target sectors within a range of -5 dB to 5 dB by adjusting the scaling of sounds both inside and outside the target sectors. The mixture of reverberated signals from speakers inside the target sectors, as captured by the reference microphone, served as our ground truth.
@@
Our complete dataset comprises 45,000 mixtures recorded across 5 different rooms. For evaluation, we employed leave-one-out validation (Wong, 2015), where we reserve one room’s data as the testing set to ensure that all test environments remain completely unseen during training process. Additionally, we allocated 1200 samples from the remaining data for validation.
Training methodology. During training, we maintained the smallest audio chunk at 8 ms (192 samples), with 4 ms lookahead. Accordingly, we configure the STFT and iSTFT window size as 288 samples with a shift length of 192. For input signals, we apply zero padding with a size of 96 at the signal end.
We apply a loss function based on the negative SI-SDR from asteriod package (Pariente et al., 2020), defined as follows:
@@
(7) $\mathcal{L}_{\text{SI-SDR}}$ =$=\begin{cases}\lambda\|\hat{\mathbf{s}}-\mathbf{s}\|_{1}&\text{nospeakers in target area}\\ -10\log\left[\frac{\|\alpha\cdot\mathbf{s}\|^{2}}{\|\alpha\cdot\mathbf{s}-\hat{\mathbf{s}}\|^{2}}\right]&\text{otherwise}\end{cases}$
Here $\mathbf{s}$ is the target clean sound, $\hat{\mathbf{s}}$ is the output sound from the network, $\|\cdot\|_{1}$ is the L1-norm, and $\lambda$ = 50 is a weighting factor, $\alpha$ can be calculated by $\alpha=\frac{\hat{\mathbf{s}}^{T}\mathbf{s}}{\|\mathbf{s}\|^{2}}$ We trained all models for 150 epochs, and apply Adam optimizer (Kingma and Ba, 2014) with an initial learning rate as 5e-3. The learning rate was increased from 5e-4 to 5e-3 over 10 epochs, held constant for 20 epochs, and then reduced by 0.95 every 2 epochs. We also applied data augmentation (Wei et al., 2020), including Time Shift, Time Stretch, Frequency Mask and Gain Perturbation during training, with a probability of 30% for each augmentation methods.
@@ 4.1. Model results
Figure 10. Stratified analysis by system variables. (a) Comparison of SI-SDR improvement across microphone arrays of varying sizes versus our SonicSieve system. (b) Average SI-SDR improvement across the 3 rooms (Room A, B, C) in our dataset, computed from 3 random locations (location 0, 1, 2) per room. Results are categorized by scenarios with 1 or 2 target speakers and averaged over 3000 samples per location.
@@
This figure presents a stratified analysis of SI-SDR improvement across different system variables, split into two bar charts: (a) Left panel compares SI-SDR improvement (y-axis, in dB ranging from 2-5.2 dB) across microphone arrays of varying sizes versus the SonicSieve system. The x-axis shows different configurations: 2-mic:  2.0 dB, 3-mic:  3.7 dB, 4-mic:  4.0 dB, 5-mic:  4.1 dB, 2-mic (SonicSieve):  4.5 dB (shown with green diagonal stripes pattern), 6-mic:  5.1 dB. This demonstrates that the 2-mic SonicSieve system achieves performance comparable to arrays with more microphones, outperforming the 5-mic arrays while using fewer hardware resources. (b) Right panel shows SI-SDR improvement (y-axis, in dB ranging from 0-6 dB) across 8 different room-device locations (A-0, A-1, A-2, B-0, B-1, B-2, C-0, C-1, C-2) on the x-axis. Each location has two bars: Blue bars: 1 Speaker scenario (values ranging from  3.8 to  5.7 dB). Red striped bars: 2 Speakers scenario (values ranging from  3.2 to  4.8 dB). In each location, the blue bar is higher than the red bar, meaning that the 1 speaker scenario is simpler than 2 speakers scenario.
Fig. 9(c) shows the relationship between the input SI-SDR of the speech samples in our dataset and the output SI-SDR after processing with our model when 1, 2, or 3 sectors are selected. The diagonal dashed line represents zero improvement, where the output SI-SDR equals the input. Points above this line indicate enhanced speech quality after processing. As evident from the figure, the vast majority of our test samples (86.6%) demonstrate positive enhancement, with output SI-SDR values consistently higher than their original values. Notably, the one-sector selection configuration (represented by red dots) shows the fewest samples (9.9%) falling below the zero-improvement line compared to two- and three-sector selections (15.2% and 16.3%, respectively), confirming our earlier observation that the task becomes more challenging as the number of selected sectors increases.
@@
Fig. 10(a) shows the SI-SDRi, computed as the difference in SI-SDR between the sound mixture at the reference microphone and the enhanced speech output from the model. As the number of microphone channels increases from 2 to 6, SI-SDRi improves from 2.1 dB to 5.1 dB, due to the increased spatial diversity. SonicSieve achieves an SI-SDRi of 4.4 dB, outperforming the 5-microphone configuration (4.1 dB), and approaching the performance of the 6-microphone configuration (5.1 dB), demonstrating the effectiveness of our whole system.
Generalization across real-world environments. The previous results were developed using datasets from Rooms B–E and evaluated in Room A. To further evaluate generalizability of SonicSieve across environments, we conducted leave-one-room-out cross-validation for Rooms B and C. Rooms D and E were excluded from this analysis due to data collection inconsistencies in the data collection procedure (with portions of data missing). Fig. 10(b) presents the SI-SDRi performance across various device locations in different rooms, with each label indicating the test room (A, B, or C) and device position (0, 1, or 2) within that room. For single-speaker scenarios (blue bars), our system achieves consistent performance across all test conditions, with SI-SDRi values ranging from 3.9 to 5.7 dB. The highest performance is observed at location A-2 (5.7 dB), while the lowest occurs at C-1 (3.9 dB). Similarly, for two-speaker scenarios (red hatched bars), our system maintains stable performance across all environments, with SI-SDRi values between 3.3 and 4.8 dB. The highest performance is observed at location A-2 (4.8 dB), while the lowest occurs at C-2 (3.3 dB).
@@
Figure 12. User interfaces for directional speech extraction. (a) Users can select the desired sector(s) using radial selectors, toggle switches or buttons on a linear bar. (b) UI evaluation setup where participants had to select one or both speakers in different scenarios. (c) Selection accuracy results for containing the target speaker. (d) Mean opinion scores for each UI. (e) Subjective confidence scores for accurately selecting target speaker(s).
[…]
This figure presents a comprehensive evaluation of user interfaces for directional speech extraction across five panels: (A) Top-left panel shows three interface types for selecting directional sectors: Radial Selectors: Fan-shaped color-coded interfaces with 6 and 9 sectors. Toggle Switches: Two rows of color-coded buttons representing different angle ranges for 6 and 9 sectors cases. Linear Bar: A horizontal bar with colored sections corresponding to different angle sectors. (B) Top-right panel shows the UI evaluation setup with two speakers (left A and right B) positioned at different locations in a room. A smartphone with the interface is visible in the foreground, with colored cone overlays indicating directional selection zones. (C) Bottom-left bar chart displays ”Likelihood of Usage” (y-axis, 1-5 scale) for six interface variations: Radial Selector (9-Sector):  4.5 (highest). Radial Selector (6-Sector):  4.8 (highest). Toggle Switch (9-Sector):  2.1 (lowest). Toggle Switch (6-Sector):  2.3 (lowest). Linear Bar (9-Sector):  3.1 (medium). Linear Bar (6-Sector):  3.4 (medium). (D) Bottom-center bar chart shows ”Selection Accuracy (percentage)” (y-axis, 0-100 percent) for the same six interfaces: Radial interfaces:  90-95 percentage. Toggle interfaces:  85-95 percent. Linear interfaces:  95-98 percent. (E) Bottom-right bar chart displays ”Confidence” scores (y-axis, 1-5 scale) across four selection restriction scenarios: Unlimited Sectors (9-Sector):  4.2. One Sector (9-Sector):  3.5. Unlimited Sectors (6-Sector):  5. One Sector (6-Sector):  4.1. The results indicate that radial selectors were most preferred by users, while linear bars achieved the highest selection accuracy. Users also expressed higher confidence when allowed to select unlimited sectors versus being restricted to one sector.

### v2

@@ 3.4. Directional speech extraction network
Feature Extraction. The core challenge in our approach is extracting the spatial cues embedded by the microstructure while remaining robust to environmental variations and multiple sound sources. While we cannot directly recover the microstructure’s direction-specific filter $M_{\theta}(f)$ due to its entanglement with unknown source signals and environmental effects, we can compute inter-channel difference features that preserve the essential directional information.
We extract interchannel phase differences (IPD) and interchannel level differences (ILD)—well-established spatial cues in binaural hearing research (Blauert, 1997; Gu and Luo, 2024):
(7) $\text{IPD}[f,t]$ $=\angle X_{2}[f,t]-\angle X_{1}[f,t]$
[…]
The IPD captures phase relationships between the two microphones, which encode timing differences created by the microstructure’s directional filtering. The ILD captures magnitude differences that reflect the microstructure’s direction-dependent attenuation and amplification patterns. Together, these features encode the spatial signatures that enable directional discrimination.
To facilitate neural network processing, we convert the circular IPD values into continuous representations using trigonometric functions:
We normalize each feature across frequency bands to ensure consistent scaling and improve training stability. For ILD, per-frequency normalization to zero mean and unit variance reduces the impact of frequency-dependent variations inherent in the microstructure design. The cosine and sine IPD components receive similar normalization to balance their influence across frequencies where phase relationships have varying reliability.
The complete feature set combines the real and imaginary STFT components from both microphone channels (4 channels) with our three spatial features, yielding 7 feature channels total. A 2D convolutional block encodes these into the representation $X\in\mathbb{R}^{c\times F\times\frac{(T+\sigma-L)}{\delta}}$, where $c$ is the expanded feature dimension.
Angle Encoder. We represent user directional selections through a 6-sector encoding system. Selected sectors receive value 1 in a one-hot vector, while adjacent sectors receive 0.25 to provide smooth spatial context. This vector undergoes embedding refinement, fully connected transformations, and reshaping to create directional embeddings $E(\mathbf{x})$ matching our feature dimensions. After LayerNorm, these embeddings condition the separation blocks through Feature-wise Linear Modulation (FiLM).
Real-time Separation Blocks. Our processing pipeline consists of $N$ sequential LSTM-based blocks optimized for real-time performance. Each block applies LayerNorm and PReLU activation, followed by convolutional downsampling to reduce computational cost for the bidirectional LSTM. After frequency-domain processing, deconvolution with skip connections restores original dimensions.
A causal LSTM with fully connected layer processes temporal relationships while maintaining real-time constraints. The FiLM layer integrates directional information by modulating features according to $E(\mathbf{x})$:
(9) $E_{sep}=X^{\prime}\odot\text{Conv1D}(E(\mathbf{x}))+\text{Conv1D}(E(\mathbf{x}))$
where $X^{\prime}$ represents processed features and $\odot$ denotes element-wise multiplication.
The final output undergoes transposed convolution and inverse STFT to produce time-domain audio. For continuous processing, we output $T$ samples for playback while retaining $\sigma$ samples for subsequent chunk processing, enabling seamless real-time operation.
@@
Figure 10. Real-world evaluation setup. (a–e) Five different test environments used for evaluating our system. The room dimensions are as follows (in meters): length $\times$ width $\times$ height. (a) Room A, 6.1 $\times$ 4.6 $\times$ 3.7; (b) Room B, 4.9 $\times$ 3.6 $\times$ 3.5; (c) Room C, 6.1 $\times$ 4.6 $\times$ 3.7; (d) Room D, 7.0 $\times$ 4.9 $\times$ 2.7; (e) Room E, 11.5 $\times$ 9.1 $\times$ 2.9. In each room we play sound samples from an external speaker (model name: JBL Flip 5), and record it with multiple devices on our data capture rig.
@@ 3.5. Real-world dataset
Motivation and Overview. Acoustic microstructures encode spatial cues through complex interactions between incoming sound waves and their geometric features. However, these spatial signatures can be significantly altered by environmental acoustic properties such as reverberation, room geometry, and background noise (Wu et al., 2016). A model trained solely on clean, simulated data may fail to generalize across the diverse acoustic environments encountered in real-world deployments.
To address this challenge, we construct a comprehensive real-world dataset that captures the acoustic variability our system will encounter in practice. Rather than relying on synthetic room impulse responses or simulated acoustics, we physically deploy our hardware in multiple real environments and record actual speech playback. This approach ensures our training data reflects the true complexity of how our microstructure interacts with real acoustic spaces, including reflections, ambient noise, and other environmental factors that synthetic approaches may not accurately model.
Our data collection process involves five distinct rooms with varying acoustic properties, multiple speaker-receiver configurations, and systematic coverage of all directional sectors. This methodology produces a diverse dataset that enables robust performance across different meeting rooms, conference spaces, and classrooms.
@@
Data Sources. We use speech content from the VCTK (Veaux et al., 2017) and LibriTTS (Zen et al., 2019) datasets, which provide high-quality recordings of human speech from multiple speakers. All audio clips are resampled to 24 kHz to balance audio quality with real-time processing requirements on mobile devices.
@@
Collection Setup and Procedure. Our data collection setup consists of a JBL Flip 5 speaker [^4] https://www.jbl.com/bluetooth-speakers/JBL+FLIP+5-.html for audio playback and a receiver rig equipped with both a MiniDSP microphone array [^5] https://www.minidsp.com/products/usb-audio-interface/uma-8-microphone-array and our SonicSieve system (Motorola Edge smartphone with microstructure). The receiver rig includes six LED lasers that divide the surrounding space into 30° sectors, enabling precise speaker placement.
We systematically collect data across five different rooms (meeting rooms, conference rooms, and classrooms) as shown in Fig. 10. For each room, we place the receiver rig at 3 randomly selected positions. At each receiver position, we place the speaker at 5 different locations distributed across all 6 directional sectors. For each speaker-receiver configuration, we play 6 random 5-second speech clips, resulting in 2,700 individual recordings (6 clips × 6 sectors × 5 speaker positions × 3 receiver positions × 5 rooms).
Mixture Generation and Dataset Statistics. To create training data for different directional extraction scenarios, we generate synthetic mixtures in software by combining recordings from different sectors. We create three types of mixtures: single-target (6 possible sector selections), two-target (15 combinations), and three-target (20 combinations) scenarios.
@@
For each room and receiver position, this produces 1,200 single-target mixtures, 900 two-target mixtures, and 900 three-target mixtures. We control the Signal-to-Noise Ratio (SNR) between target and interfering sectors within -5 to 5 dB by scaling the audio appropriately. The ground truth for each mixture is the clean combination of target sector signals as recorded by the reference microphone.
@@
Our complete dataset comprises 45,000 mixtures across 5 rooms. We employ leave-one-room-out validation (Wong, 2015) to ensure generalization, reserving one room’s data (9,000 mixtures) for testing while using the remaining rooms for training. We additionally reserve 1,200 samples from the training data for validation.
Training Methodology. We process audio in 8 ms chunks (192 samples at 24 kHz) with 4 ms lookahead for real-time performance. The STFT uses a window size of 288 samples with 192-sample shifts, and we apply 96-sample zero padding at signal boundaries.
We optimize using a Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) loss from the Asteroid package (Pariente et al., 2020):
@@
(10) $\mathcal{L}_{\text{SI-SDR}}$ $=\begin{cases}\lambda\|\hat{\mathbf{s}}-\mathbf{s}\|_{1}&\text{no speakers in target area}\\ -10\log\left[\frac{\|\alpha\cdot\mathbf{s}\|^{2}}{\|\alpha\cdot\mathbf{s}-\hat{\mathbf{s}}\|^{2}}\right]&\text{otherwise}\end{cases}$
where $\mathbf{s}$ is the target clean sound, $\hat{\mathbf{s}}$ is the network output, $\lambda=50$ is a weighting factor, and $\alpha=\frac{\hat{\mathbf{s}}^{T}\mathbf{s}}{\|\mathbf{s}\|^{2}}$ provides scale invariance.
We train for 150 epochs using the Adam optimizer (Kingma and Ba, 2014). The learning rate starts at $5\texttimes 10^{-4}$, increases to $5\texttimes 10^{-3}$ over 10 epochs, remains constant for 20 epochs, then decays by 0.95 every 2 epochs. We apply data augmentation (Time Shift, Time Stretch, Frequency Mask, and Gain Perturbation) with 30% probability for each method (Wei et al., 2020).
@@ 4.1. Model results
Figure 11. System performance compared to the baselines without microstructure (a) Effect of selected sector index on SI-SDR improvement (b) SI-SDR improvement as a function of the number of simultaneously selected sectors.
[…]
This figure presents three graphs evaluating system performance through SI-SDR (Scale-Invariant Signal-to-Distortion Ratio) measurements across different conditions: (a) A line graph showing the effect of selected sector index (x-axis, numbered 1-6) on SI-SDR improvement in dB (y-axis, ranging 2-7 dB). Two lines compare performance: With microstructure (light green line with circles): Shows significantly higher SI-SDR values (between 3.6-6.8 dB) with a U-shaped pattern where sectors 1 and 6 perform best. The SI-SDR number for each sector is 6.4, 4.7, 4.4, 4.1, 3.6, 6.8. Without microstructure (dark blue line with squares): Shows consistently lower performance (between 1.9-2.8 dB) with a slight downward trend as sector index increases. The SI-SDR number for each sector is 2.7, 2.0, 2.5, 2.4, 2.1, 1.9. (b) A bar graph comparing SI-SDR improvement (y-axis, in dB) across different numbers of simultaneously queried sectors (x-axis, 1-3 sectors). For each number of sectors, two bars compare: With microstructure (green bars): Shows 5.0 dB for 1 sector, 4.2 dB for 2 sectors, and 3.9 dB for 3 sectors. Without microstructure (blue bars): Shows substantially lower values of 2.3 dB, 1.9 dB, and 1.7 dB respectively.
@@
Fig. 12(a) shows the relationship between the input SI-SDR of the speech samples in our dataset and the output SI-SDR after processing with our model when 1, 2, or 3 sectors are selected. The diagonal dashed line represents zero improvement, where the output SI-SDR equals the input. Points above this line indicate enhanced speech quality after processing. As evident from the figure, the vast majority of our test samples (86.6%) demonstrate positive enhancement, with output SI-SDR values consistently higher than their original values. Notably, the one-sector selection configuration (represented by red dots) shows the fewest samples (9.9%) falling below the zero-improvement line compared to two- and three-sector selections (15.2% and 16.3%, respectively), confirming our earlier observation that the task becomes more challenging as the number of selected sectors increases.
Figure 12. Overall system performance. (a) Input vs. output SI-SDR of our model across all samples, grouped by 1-sector, 2-sector, and 3-sector selections. Points above the dashed line indicate cases where SI-SDR improved. (b) Comparison of SI-SDR improvement across microphone arrays of varying sizes versus our SonicSieve system.
@@
(a) A scatter plot showing input vs. output SI-SDR (x-axis: Original SI-SDR ranging from -25 to 15 dB; y-axis: Output SI-SDR ranging from -30 to 20 dB). Points are color-coded by query type: 1 Sector Query (red dots), 2 Sectors Query (blue dots), 3 Sectors Query (orange dots). A diagonal dashed line indicates ”Zero improvement” reference. Most data points cluster above this line (especially in the -5 to 5 dB original SI-SDR range), demonstrating positive enhancement, with the upper-right quadrant labeled ”Enhancement” in green. (b) A bar plot compares SI-SDR improvement (y-axis, in dB ranging from 2-5.2 dB) across microphone arrays of varying sizes versus the SonicSieve system. The x-axis shows different configurations: 2-mic:  2.0 dB, 3-mic:  3.7 dB, 4-mic:  4.0 dB, 5-mic:  4.1 dB, 2-mic (SonicSieve):  4.5 dB (shown with green diagonal stripes pattern), 6-mic:  5.1 dB. This demonstrates that the 2-mic SonicSieve system achieves performance comparable to arrays with more microphones, outperforming the 5-mic arrays while using fewer hardware resources.
@@
Fig. 12(b) shows the SI-SDRi, computed as the difference in SI-SDR between the sound mixture at the reference microphone and the enhanced speech output from the model. As the number of microphone channels increases from 2 to 6, SI-SDRi improves from 2.1 dB to 5.1 dB, due to the increased spatial diversity. SonicSieve achieves an SI-SDRi of 4.4 dB, outperforming the 5-microphone configuration (4.1 dB), and approaching the performance of the 6-microphone configuration (5.1 dB), demonstrating the effectiveness of our whole system.
Generalization across real-world environments. The previous results were developed using datasets from Rooms B–E and evaluated in Room A. To further evaluate generalizability of SonicSieve across environments, we conducted leave-one-room-out cross-validation for Rooms B and C. Rooms D and E were excluded from this analysis due to data collection inconsistencies in the data collection procedure (with portions of data missing). Fig. 13(a) presents the SI-SDRi performance across various device locations in different rooms, with each label indicating the test room (A, B, or C) and device position (0, 1, or 2) within that room. For single-speaker scenarios (blue bars), our system achieves consistent performance across all test conditions, with SI-SDRi values ranging from 3.9 to 5.7 dB. The highest performance is observed at location A-2 (5.7 dB), while the lowest occurs at C-1 (3.9 dB). Similarly, for two-speaker scenarios (red hatched bars), our system maintains stable performance across all environments, with SI-SDRi values between 3.3 and 4.8 dB. The highest performance is observed at location A-2 (4.8 dB), while the lowest occurs at C-2 (3.3 dB).
Impact of smartphone thickness. We evaluated how smartphone thickness affects system performance since our system is designed to generalize to different smartphone geometries. Using our default prototype with the Motorola Edge 2024 (10 mm thickness), we modeled the effects of thicker devices by adding foam tape between the phone screen and microstructure to simulate 12 mm and 14 mm thickness. Fig. 13(b) shows that SonicSieve maintains robust performance across this thickness range. For single-speaker scenarios, SI-SDRi decreases slightly from 4.84 dB at 10 mm to 4.79 dB at 12 mm and 4.43 dB at 14 mm. Similarly, for two-speaker scenarios, performance drops from 3.77 dB at 10 mm to 3.65 dB at 12 mm and 3.21 dB at 14 mm. While thicker devices show modest performance reduction, the degradation is relatively small (less than 0.6 dB across the tested range). Since most off-the-shelf smartphones have thickness between 8–12 mm, our system demonstrates good compatibility with existing device form factors.
Figure 13. Stratified analysis by system variables. (a) Average SI-SDR improvement across the 3 rooms (Room A, B, C) in our dataset, computed from 3 random locations (location 0, 1, 2) per room. Results are categorized by scenarios with 1 or 2 target speakers and averaged over 3000 samples per location. (b) Average SI-SDR improvement with different smartphone thickness. The experiment is conducted at the B-0 location with three prototypes.
@@
(a) This figure presents a stratified analysis of SI-SDR improvement across different system variables, split into two bar charts: (a) Right panel shows SI-SDR improvement (y-axis, in dB ranging from 0-6 dB) across 8 different room-device locations (A-0, A-1, A-2, B-0, B-1, B-2, C-0, C-1, C-2) on the x-axis. Each location has two bars: Blue bars: 1 Speaker scenario (values ranging from  3.8 to  5.7 dB). Red striped bars: 2 Speakers scenario (values ranging from  3.2 to  4.8 dB). In each location, the blue bar is higher than the red bar, meaning that the 1 speaker scenario is simpler than 2 speakers scenario. (b) is a bar chart showing SI-SDR improvement (SI-SDRi) in decibels on the y-axis, ranging from 0 to 6 dB, versus smartphone thickness on the x-axis with three categories: 10 mm, 12 mm, and 14 mm. For each thickness category, there are two bars: a solid blue bar representing ”1 Speaker” scenarios and a red hatched bar representing ”2 Speakers” scenarios. The blue bars show values of approximately 4.8 dB for 10 mm, 4.8 dB for 12 mm, and 4.4 dB for 14 mm thickness. The red hatched bars show consistently lower values of approximately 3.8 dB for 10 mm, 3.7 dB for 12 mm, and 3.2 dB for 14 mm thickness. Both single-speaker and two-speaker performance show a gradual decline as smartphone thickness increases, with single-speaker scenarios consistently outperforming two-speaker scenarios across all thickness values. A legend in the upper right corner identifies the blue bars as ”1 Speaker” and red hatched bars as ”2 Speakers.”